# TinyGraphEstimator: Adapting Lightweight Language Models for Graph Structure Inference

### v1

@@ Abstract
Graphs provide a universal framework for representing complex relational systems, and inferring their structural properties is a core challenge in graph analysis and reasoning. While large language models have recently demonstrated emerging abilities to perform symbolic and numerical reasoning, the potential of smaller, resource-efficient models in this context remains largely unexplored. This paper investigates whether compact transformer-based language models can infer graph-theoretic parameters directly from textual graph representations. To enable systematic evaluation, we introduce the TinyGraphEstimator dataset - a balanced collection of connected graphs generated from multiple random graph models and annotated with detailed structural metadata. We evaluate several small open models on their ability to predict key graph parameters such as density, clustering, and chromatic number. Furthermore, we apply lightweight fine-tuning using the Low-Rank Adaptation (LoRA) technique, achieving consistent improvements across all evaluated metrics. The results demonstrate that small language models possess non-trivial reasoning capacity over graph-structured data and can be effectively adapted for structural inference tasks through efficient parameter tuning.
@@ 1 INTRODUCTION
To address this gap, we investigate whether small transformer-based language models can infer structural graph parameters directly from serialized graph representations. The models operate on deterministic graph formats - edge lists or adjacency lists - and are tasked with predicting graph-theoretic measures from these structured inputs. This setup allows us to examine whether compact language models can generalize reasoning patterns over relational structures without relying on large-scale model capacity or specialized graph algorithms.
@@ 2 RELATED WORK
Recent advances in large language models (LLMs) have demonstrated impressive performance in tasks requiring symbolic reasoning, numerical computation, and pattern generalization [Brown et al., 2020, Bubeck et al., 2023]. Studies have shown that transformer-based architectures can internalize relational structures [Wang et al., 2025] and perform graph-related reasoning when provided with suitable prompts or structured inputs. Nevertheless, most prior work focuses on large-scale models with hundreds of billions of parameters, whose reasoning capabilities emerge from scale rather than explicit adaptation. In contrast, we investigate whether small language models - under 5B parameters - can infer precise graph-theoretic quantities such as node count, triangle number, or chromatic number. This shifts the focus from emergent reasoning in massive models to lightweight, interpretable reasoning in compact architectures.
@@ 3 PROPOSED SOLUTION
The proposed approach evaluates and enhances the ability of small language models to infer a wide range of graph-theoretic parameters directly from structured graph representations. Each model is prompted to predict a structured set of key structural properties that collectively describe its topology and complexity, based on the graph’s edge list or adjacency list representation. The inferred parameters include the number of nodes, edges, and density, as well as detailed degree-based statistics such as minimum degree, mean degree, maximum degree, and degree standard deviation. In addition to these fundamental attributes, we evaluate the models’ ability to estimate higher-order structural metrics including the total number of triangles, average clustering coefficient, transitivity, average shortest path length, diameter, chromatic number, and global efficiency. Collectively, these parameters capture both local and global aspects of graph organization, providing a comprehensive benchmark for assessing reasoning over graph structures.
@@ 3.1 Models Selection
To investigate the ability of small language models to infer structural properties of graphs, we selected a group of compact yet high-performing models that combine efficient deployment with strong reasoning capacity. Our goal was to identify models that are powerful enough to capture graph-theoretic patterns while remaining computationally accessible for systematic experimentation. Model selection was guided by comparative performance metrics reported on the LLM-Stats leaderboard [LLM Stats, 2024], focusing on models with fewer than 5 billion parameters that demonstrate competitive reasoning accuracy across diverse benchmarks. In particular, we prioritized models offering extended context lengths of at least 128,000 tokens, a crucial requirement for processing full graph representations.
The final selection includes Qwen-2.5-3B [Yang et al., 2025], Llama 3.2-3B [Grattafiori et al., 2024], and Phi-4-mini-4B [Abdin et al., 2024] (all in instruct variant). These models were chosen because they consistently rank among the strongest sub-5B architectures in reasoning and comprehension tasks while maintaining full open availability for reproducible research. Their extended context capacity allows complete representations of graphs to be processed in a single inference window. Moreover, their compact architecture enables efficient fine-tuning and evaluation on standard hardware, making them ideal candidates for exploring how small-scale language models internalize and generalize graph-structural information.
@@ 3.2 Dataset Construction
To evaluate the ability of small language models to infer graph-theoretic properties, we constructed a balanced synthetic dataset of connected graphs generated from three canonical random graph models: Erdős–Rényi (ER), Barabási–Albert (BA), and Watts–Strogatz (WS). The ER model produces graphs where each possible edge occurs independently with a fixed probability, capturing purely random connectivity [Erdös and Rényi, 1959]. The BA model generates scale-free networks through preferential attachment, emphasizing hub formation and heavy-tailed degree distributions [Barabási and Albert, 1999]. The WS model interpolates between regular lattices and random graphs, combining high clustering with short path lengths [Watts and Strogatz, 1998].
@@
Graphs are stored in multiple formats (edgelist, adjacency list, and JSON node-link representation), accompanied by both compact and full metadata that capture key structural properties including node and edge counts, density, degree statistics, clustering coefficient, transitivity, chromatic number, diameter, and triangle count. The dataset consists of 1,000 graphs for training and 100 graphs for testing, evenly distributed across the four models. This design provides a well-controlled yet structurally varied benchmark for assessing the generalization and reasoning abilities of lightweight language models in inferring global and local graph parameters.
@@ 3.3 Training Setup
We fine-tune each base model with parameter-efficient supervised instruction tuning using the TRL SFTTrainer [von Werra et al., 2020] and LoRA adapters. Training data consist of paired promptcompletion examples: the prompt encodes the normalized graph (edge or adjacency list), the fixed target schema, and concise instructions; the completion is the gold JSON with exactly fourteen fields. To avoid learning spurious prompt tokens and to sharpen output formatting, we use completion-only loss. LoRA adapters are applied to attention and MLP projection modules with rank r=16r{=}16, $\alpha{=}32$, and dropout $0.05$, leaving the base weights frozen. We train with sequence length $2048$, batch size $2$ and gradient accumulation $8$, cosine learning-rate schedule with warmup ratio $0.03$, learning rate $2\!\times\!10^{-4}$, and one epoch over the training set. Mixed precision uses bfloat16 when available.
@@
All experiments (training and validation) were executed on a single NVIDIA RTX 3090 GPU (Ampere, 24 GB GDDR6X, Tensor Cores) in a standard Linux x86_64 environment.
@@ 3.4 Validation Protocol
We evaluate model outputs with a schema- and tolerance-aware validation pipeline that mirrors the training prompt format while remaining model-agnostic. For each sample, the script loads the graph data and gold metadata, detects whether the input is an edge list or adjacency list, and normalizes it to a canonical textual form. A structured prompt then instructs the model to return only a JSON object matching a fixed schema of given fields (nodes, edges, density, degree statistics, triangles, clustering, transitivity, path lengths, diameter, chromatic number, global efficiency). Inference is run on a base model or a base model augmented with LoRA adapters (if provided), with optional JSON-constrained decoding via lm-format-enforcer [Gat, 2024] to prevent malformed outputs. The first JSON object in the response is extracted and strictly checked against the schema (type enforcement with graceful acceptance of null when a value is not produced). Accuracy is computed per field: integers require exact equality (after integer casting), while real-valued fields use absolute and relative tolerances to decide correctness and record signed errors.
Table 1: Normalized Root Mean Squared Error (NRMSE) for graph parameter inference across small language models. Each model is evaluated in plain (zero-shot) and fine-tuned variants. Values are normalized by the range of each parameter and averaged over the test set.
Parameter Llama-3.2-3B Qwen-2.5-3B Phi-4-mini-4B
Plain Fine-tuned Plain Fine-tuned Plain Fine-tuned
Nodes 2.36 0.51 2.36 0.37 2.36 0.38
Edges 0.54 0.18 0.54 0.33 0.54 0.27
Density 0.69 0.24 0.58 0.22 0.64 0.32
Degree Min 0.43 0.23 0.43 0.35 0.43 0.23
Degree Mean 0.63 0.22 0.63 0.32 0.63 0.33
Degree Max 0.58 0.26 0.58 0.19 0.58 0.35
Degree Std 0.52 0.37 0.54 0.27 0.54 0.41
Triangles Total 0.32 0.23 0.32 0.29 0.32 0.30
Average Clustering 0.54 0.20 0.54 0.19 0.54 0.23
Transitivity 0.54 0.24 0.55 0.23 0.55 0.24
Average Shortest Path Length 0.35 0.14 0.35 0.09 0.41 0.23
Diameter 0.28 0.13 0.29 0.08 0.29 0.22
Chromatic Number 0.69 0.19 0.55 0.25 0.70 0.30
Global Efficiency 1.01 0.21 1.37 0.16 1.06 0.31
@@ 4 RESULTS
The experimental results, summarized in Table 1, highlight consistent trends in the ability of small language models to infer graph-theoretic properties. In the zero-shot setting, all models perform poorly across most parameters, indicating limited inherent capability for structural reasoning. While simple counts are occasionally inferred correctly, higher-order statistics - particularly those involving multi-node dependencies or global structure, such as triangles, clustering coefficients, path lengths, and chromatic number - show substantial deviations from the ground truth. This observation confirms that, without explicit adaptation, compact models lack the internal procedural knowledge required for accurate structural reasoning. Quantitatively, both the mean absolute error (MAE) and root mean squared error (RMSE) are high across almost all metrics, reflecting limited numerical precision and inconsistency between samples.
After parameter-efficient fine-tuning using the LoRA approach, performance improves consistently for every evaluated field. The reductions in MAE and RMSE across model families demonstrate that even lightweight adaptation can significantly enhance reasoning over graph-structured data. Improvements are most pronounced for mid-level structural descriptors - degree statistics, density, and clustering-related metrics - indicating that fine-tuning enables models to internalize relational patterns and approximate algorithmic reasoning without direct supervision over computation steps. More complex properties such as average shortest path length and chromatic number remain challenging but also show measurable gains. Overall, the adapted models produce more stable, schema-consistent, and numerically coherent predictions, suggesting that small transformers can, when properly guided, develop an emergent understanding of graph topology sufficient for practical inference tasks. These results collectively validate the effectiveness of the TinyGraphEstimator dataset and the LoRA fine-tuning strategy in unlocking latent reasoning capacity within compact language models.
@@ 5 LIMITATIONS
While this study demonstrates the potential of small language models for graph-structural reasoning, several limitations should be acknowledged. First, although LoRA fine-tuning consistently improves inference accuracy, the models still exhibit challenges in predicting complex or global parameters such as chromatic number and average shortest path length, which require deeper combinatorial reasoning. Second, the evaluation is based on synthetic graphs generated from a selected set of well-established random graph models, which provide controlled variability but may not reflect the full diversity of real-world network structures. Incorporating additional graph types or domain-specific datasets in future work could further strengthen the assessment of model generalization. Third, the experiments rely on serialized graph formats - edge lists or adjacency lists - that, while simple and interpretable, may not capture complex or densely connected structures as effectively as more expressive representations. Future work could explore alternative input forms or hybrid architectures that integrate symbolic and neural reasoning. Despite these limitations, the results highlight promising directions for lightweight, adaptable language models in graph-theoretic inference.

### v2

@@ Abstract
Graphs provide a universal framework for representing complex relational systems, and inferring their structural properties is a core challenge in graph analysis and reasoning. While large language models have recently demonstrated emerging abilities to perform symbolic and numerical reasoning, the potential of smaller, resource-efficient models in this context remains largely unexplored. This paper investigates whether compact transformer-based language models can infer graph-theoretic parameters directly from graph representations. To enable systematic evaluation, we introduce the TinyGraphEstimator dataset - a balanced collection of connected graphs generated from multiple random graph models and annotated with detailed structural metadata. We evaluate several small open models on their ability to predict key graph parameters such as density, clustering, and chromatic number. Furthermore, we apply lightweight fine-tuning using the Low-Rank Adaptation (LoRA) technique, achieving consistent improvements across all evaluated metrics. The results demonstrate that small language models possess non-trivial reasoning capacity over graph-structured data and can be effectively adapted for structural inference tasks through efficient parameter tuning.
@@ 1 INTRODUCTION
To address this gap, we investigate whether small transformer-based language models can infer structural graph parameters directly from serialized graph representations. The models operate on deterministic graph format - edge list - and are tasked with predicting graph-theoretic measures from these structured inputs. This setup allows us to examine whether compact language models can generalize reasoning patterns over relational structures without relying on large-scale model capacity or specialized graph algorithms.
@@ 2 RELATED WORK
Recent advances in large language models (LLMs) have demonstrated impressive performance in tasks requiring symbolic reasoning, numerical computation, and pattern generalization [Brown et al., 2020, Bubeck et al., 2023]. Studies have shown that transformer-based architectures can internalize relational structures [Wang et al., 2025] and perform graph-related reasoning when provided with suitable prompts or structured inputs. Nevertheless, most prior work focuses on large-scale models with hundreds of billions of parameters, whose reasoning capabilities emerge from scale rather than explicit adaptation. In contrast, we investigate whether small language models - under 4B parameters - can infer precise graph-theoretic quantities. This shifts the focus from emergent reasoning in massive models to lightweight, interpretable reasoning in compact architectures.
@@ 3 PROPOSED SOLUTION
The proposed approach evaluates and enhances the ability of small language models to infer a wide range of graph-theoretic parameters directly from structured graph representations. Each model is prompted to predict a structured set of key structural properties that collectively describe its topology and complexity, based on the graph’s edge list representation. The inferred parameters include the detailed degree-based statistics such as density, minimum degree, mean degree, maximum degree, and degree standard deviation. In addition to these fundamental attributes, we evaluate the models’ ability to estimate higher-order structural metrics including the total number of triangles, average clustering coefficient, transitivity, average shortest path length, diameter, chromatic number, and global efficiency. Collectively, these parameters capture both local and global aspects of graph organization, providing a comprehensive benchmark for assessing reasoning over graph structures.
@@ 3.1 Models Selection
To investigate the ability of small language models to infer structural properties of graphs, we selected a group of compact yet high-performing models that combine efficient deployment with strong reasoning capacity. Our goal was to identify models that are powerful enough to capture graph-theoretic patterns while remaining computationally accessible for systematic experimentation. Model selection was guided by comparative performance metrics reported on the LLM-Stats leaderboard [LLM Stats, 2024], focusing on models with fewer than 4B parameters that demonstrate competitive reasoning accuracy across diverse benchmarks.
The final selection includes Qwen-2.5-3B [Yang et al., 2025], Llama-3.2-3B [Grattafiori et al., 2024], and Phi-4-mini (4B) [Abdin et al., 2024], all used in their Instruct variants. These models were chosen because they consistently rank among the strongest sub-4B architectures in reasoning and comprehension tasks while maintaining full open availability for reproducible research. Their extended context capacity allows complete representations of graphs to be processed in a single inference window. Moreover, their compact architecture enables efficient fine-tuning and evaluation on standard hardware, making them ideal candidates for exploring how small-scale language models internalize and generalize graph-structural information.
@@ 3.2 Dataset Construction
To the best of our knowledge, no existing dataset directly supports the task of mapping graph structures to their quantitative parameters. Therefore, we constructed a balanced synthetic dataset of connected graphs generated from three canonical random graph models: Erdős–Rényi (ER), Barabási–Albert (BA), and Watts–Strogatz (WS). The ER model produces graphs where each possible edge occurs independently with a fixed probability, capturing purely random connectivity [Erdös and Rényi, 1959]. The BA model generates scale-free networks through preferential attachment, emphasizing hub formation and heavy-tailed degree distributions [Barabási and Albert, 1999]. The WS model interpolates between regular lattices and random graphs, combining high clustering with short path lengths [Watts and Strogatz, 1998].
@@
Graphs are stored accompanied by metadata that capture key structural properties. The dataset consists of 1,200 graphs for training and 120 graphs for testing, evenly distributed across the three models. This design provides a well-controlled yet structurally varied benchmark for assessing the generalization and reasoning abilities of lightweight language models in inferring global and local graph parameters.
@@ 3.3 Training Setup
We fine-tune each base model with parameter-efficient supervised instruction tuning using the TRL SFTTrainer [von Werra et al., 2020] and LoRA adapters. Training data consist of paired promptcompletion examples: the prompt includes the normalized graph (edge list), the fixed target schema, and concise instructions; the completion is the gold JSON with expected fields. To avoid learning spurious prompt tokens and to sharpen output formatting, we use completion-only loss. LoRA adapters are applied to attention and MLP projection modules with rank r=32r{=}32, $\alpha{=}32$, and dropout $0.05$. We train with sequence length $2048$, batch size $2$ and gradient accumulation $8$, cosine learning-rate schedule with warmup ratio $0.03$, learning rate $2\!\times\!10^{-4}$, and ten epochs over the training set.
@@
All experiments (training and validation) were executed on a single NVIDIA RTX 3090 GPU (Ampere, 24GB VRAM) in a standard Linux x86_64 environment.
@@ 3.4 Validation Protocol
We evaluate model outputs with a schema- and tolerance-aware validation pipeline that mirrors the training prompt format while remaining model-agnostic. For each sample, the script loads the graph data and gold metadata and normalizes it to a canonical textual form. A structured prompt then instructs the model to return only a JSON object matching a fixed schema of given fields (nodes, edges, density, degree statistics, triangles, clustering, transitivity, path lengths, diameter, chromatic number, global efficiency). Inference is run on a base model or a base model augmented with LoRA adapters (if provided), with JSON-constrained decoding via lm-format-enforcer [Gat, 2024] to prevent malformed outputs. The first JSON object in the response is extracted and strictly checked against the schema.
3.5 Evaluation Metrics
Performance was evaluated using four standard regression metrics: the normalized root mean squared error based on range ($\mathrm{NRMSE}_{\text{range}}$), the symmetric mean absolute percentage error (sMAPE), the $R^{2}$ Accuracy, and the NRMSE Accuracy.
$\mathrm{NRMSE}_{\text{range}}$ represents the root mean squared error normalized by the range of the target variable, providing a scale-independent measure of average deviation between predicted and true values. sMAPE quantifies the mean relative difference between predictions and ground truth, computed symmetrically with respect to both magnitudes to balance over- and underestimation. $R^{2}$ Accuracy expresses the proportion of variance in the target data explained by the model, calculated as $100\times\max(0,R^{2})$, while NRMSE Accuracy converts normalized RMSE with respect to the target standard deviation into a percentage scale using $100\times(1-\mathrm{NRMSE}_{\text{std}})$. All accuracy metrics and sMAPE are reported in percentage form for comparability across parameters. Lower $\mathrm{NRMSE}_{\text{range}}$ and sMAPE values and higher accuracy scores indicate better predictive performance.
@@ 4 RESULTS
Table 1: Comparison of $\mathrm{NRMSE}_{\text{range}}$ and sMAPE (%) for graph parameter inference. Each model is evaluated in plain (zero-shot) and fine-tuned variants. Lower values indicate smaller relative errors and therefore better predictions. Fine-tuned results are underlined, and the best result for each parameter is shown in bold.
Parameter $\mathbf{NRMSE}_{\text{range}}$ sMAPE (%)
Llama-3.2-3B Qwen-2.5-3B Phi-4-mini (4B) Llama-3.2-3B Qwen-2.5-3B Phi-4-mini (4B)
Plain Fine-tuned Plain Fine-tuned Plain Fine-tuned Plain Fine-tuned Plain Fine-tuned Plain Fine-tuned
Density 0.902 0.000 0.317 0.005 0.998 0.014 72.764 0.000 56.069 0.201 92.402 0.654
Degree Min 0.349 0.100 0.280 0.090 0.270 0.088 73.697 16.717 46.771 17.376 52.291 17.393
Degree Mean 0.397 0.001 0.311 0.000 0.160 0.004 65.803 0.011 45.800 0.000 23.579 0.217
Degree Max 0.284 0.033 0.250 0.034 0.331 0.037 44.846 3.335 34.764 3.594 60.272 4.559
Degree Std 0.268 0.056 0.293 0.054 0.257 0.053 48.499 9.896 44.476 9.848 45.682 10.266
Triangles Total 0.314 0.030 0.198 0.023 0.291 0.037 163.371 14.027 86.373 9.808 177.232 16.806
Average Clustering 0.594 0.076 0.370 0.056 0.510 0.083 168.042 14.209 105.227 10.034 178.104 15.928
Transitivity 0.619 0.053 0.368 0.042 0.518 0.053 168.327 10.312 102.598 8.151 178.718 11.463
Average Shortest Path Length 0.237 0.048 0.222 0.051 0.345 0.058 26.490 2.720 32.143 2.794 134.941 3.286
Diameter 0.215 0.068 0.236 0.061 0.218 0.074 38.764 7.399 39.536 5.445 46.950 6.565
Chromatic Number 0.224 0.078 0.195 0.076 0.348 0.075 26.137 5.887 21.712 5.298 41.875 5.413
Global Efficiency 0.894 0.017 0.353 0.019 0.994 0.026 140.083 1.290 32.294 1.272 176.948 1.778
Overall (Mean) 0.441 0.047 0.283 0.043 0.437 0.050 86.402 7.150 53.980 6.152 100.750 7.861
Table 2: Comparison of models on $R^{2}$ Accuracy (%) and NRMSE Accuracy (%) across graph parameters. Higher values indicate better predictive performance and closer agreement between predicted and true graph-structural properties. Best results for each parameter are shown in bold.
Parameter R2 Accuracy (%) NRMSE Accuracy (%)
Llama-3.2-3B Qwen-2.5-3B Phi-4-mini (4B) Llama-3.2-3B Qwen-2.5-3B Phi-4-mini (4B)
Density 100.000 99.957 99.621 100.000 97.926 93.840
Degree Min 85.213 87.957 88.415 61.546 65.297 65.963
Degree Mean 99.999 100.000 99.974 99.697 100.000 98.394
Degree Max 97.924 97.885 97.415 85.591 85.456 83.921
Degree Std 95.286 95.599 95.775 78.289 79.021 79.445
Triangles Total 98.284 98.996 97.474 86.898 89.981 84.107
Average Clustering 92.539 95.845 90.945 72.686 79.617 69.908
Transitivity 96.465 97.753 96.384 81.198 85.009 80.984
Average Shortest Path Length 94.249 93.636 91.603 76.020 74.773 71.022
Diameter 88.521 90.875 86.226 66.120 69.792 62.886
Chromatic Number 86.986 87.487 87.988 63.926 64.626 65.341
Global Efficiency 99.485 99.317 98.766 92.822 91.736 88.893
Overall (Mean) 94.579 95.442 94.216 80.339 81.936 78.725
A comparison of the performance of base and fine-tuned models is presented in Table 1 (accuracy results for untrained models are near zero and therefore omitted). The results show a clear and systematic improvement across all evaluated metrics. Fine-tuning leads to a substantial reduction in normalized errors, with each model achieving markedly higher precision after adaptation. For $\mathrm{NRMSE}_{\text{range}}$, the mean error decreases from 0.441 to 0.047 for Llama-3.2-3B, from 0.283 to 0.043 for Qwen-2.5-3B, and from 0.437 to 0.050 for Phi-4-mini (4B), corresponding to an average improvement of approximately 0.340. A similar trend is observed for sMAPE, where mean values drop from 86.4% to 7.2% for Llama, from 54.0% to 6.2% for Qwen, and from 100.8% to 7.9% for Phi - an overall reduction exceeding 70 percentage points. These results demonstrate that LoRA fine-tuning consistently and substantially enhances model precision, reducing both absolute and proportional errors by more than an order of magnitude.
Table 2 summarizes the accuracy of the fine-tuned models. The results remain consistently very high, with $R^{2}$ Accuracy exceeding 94% and NRMSE Accuracy surpassing 78% across all models and parameters. Among them, Qwen-2.5-3B achieved the best overall performance, reaching a mean $R^{2}$ Accuracy of 95.4% and an NRMSE Accuracy of 81.9%. Llama-3.2-3B followed closely with 94.6% and 80.3%, while Phi-4-mini (4B) achieved 94.2% and 78.7%, respectively. Although the differences are modest, Qwen shows a consistent advantage on metrics capturing higher-order structural organization, such as transitivity, clustering, and triangle counts - properties that depend on multi-node relationships and global connectivity. This suggests that Qwen’s architecture or pretraining corpus may better encode relational and hierarchical dependencies in serialized graph input.
Across all models, accuracy remains stable across different graph types, from sparse to dense topologies, indicating robust generalization and resilience to variation in degree distribution and connectivity. In summary, fine-tuning enables small, LoRA-adapted language models to perform precise quantitative reasoning over graph-structured data, serving as reliable, high-precision estimators of graph-theoretic properties.
5 DISCUSSION
The results demonstrate that compact transformer models, when fine-tuned with parameter-efficient techniques such as LoRA, are capable of accurate and interpretable reasoning over graph-structured data. The observed improvements across all graph parameters confirm that small language models can internalize both local and global structural dependencies without requiring explicit graph neural architectures. This finding has several broader implications.
First, it suggests that structural reasoning is not an emergent property of scale alone, but can instead be induced through targeted adaptation. This challenges the assumption that only very large language models possess strong symbolic or numerical inference abilities, and indicates that efficiency-oriented architectures can deliver competitive performance under proper supervision. Second, the TinyGraphEstimator dataset and methodology provide a controlled framework for evaluating how language models interpret formal structures, which could extend beyond graphs to domains such as program analysis, knowledge graphs, or symbolic mathematics. Finally, the results point toward practical opportunities for deploying small LMs in low-resource or embedded environments, where traditional large-scale LLMs are computationally infeasible.
In essence, the strong predictive performance and generalization of fine-tuned models highlight a promising direction for bridging discrete reasoning and natural language modeling. By demonstrating that small transformers can perform accurate structural inference, this work lays the groundwork for future exploration of hybrid symbolic–neural systems and for extending lightweight reasoning models to real-world relational data.
@@ 6 LIMITATIONS
While this study demonstrates the potential of small language models for graph-structural reasoning, several limitations should be acknowledged. First, although LoRA fine-tuning consistently improves inference accuracy, the models still exhibit challenges in predicting complex or global parameters such as chromatic number and average shortest path length, which require deeper combinatorial reasoning. Second, the evaluation is based on synthetic graphs generated from a selected set of well-established random graph models, which provide controlled variability but may not reflect the full diversity of real-world network structures. Incorporating additional graph types or domain-specific datasets in future work could further strengthen the assessment of model generalization. Third, the experiments rely on serialized graph formats that, while simple and interpretable, may not capture complex or densely connected structures as effectively as more expressive representations. Future work could explore alternative input forms or hybrid architectures that integrate symbolic and neural reasoning. Despite these limitations, the results highlight promising directions for lightweight, adaptable language models in graph-theoretic inference.